(y+2)^2 = 20
solve for y my nephew and i both keep getting different answers
: (y+2)^2 = 20
solve for y my nephew and i both keep getting different answers

Answer by jim_thompson5910(9390)


(y+2)^2=20 Start with the given equation.


x+2=0+-sqrt(20) Take the square root of both sides.


y+2=sqrt(20) or y+2=-sqrt(20) Break up the "plus/minus" to form two equations.


y+2=2*sqrt(5) or y+2=-2*sqrt(5) Simplify the square root.


y=-2+2*sqrt(5) or y=-2-2*sqrt(5) Subtract 2 from both sides.


--------------------------------------


Answer:


So the solutions are y=-2+2*sqrt(5) or y=-2-2*sqrt(5).


which approximate to y=2.472 or y=-6.472